simplify 2/3 if y=6
1
Expert's answer
2019-10-31T06:22:59-0400

Solution:


Given ,

"\frac {2y} {3}"



Now, To simplify, we can plug the value y = 6 in the expression

"\frac {2y} {3}"

So,

"\frac {2y} {3} = \frac {2 (6) } {3} = \frac {12} {3} = 4"

Answer:

"\frac {2y} {3} = 4"